# INDRA Collaboration

### v1

@@ II Experimental details
Systems $E_{beam}$ $K^{(CM)}_{tot}$ $\beta^{(lab)}_{CM}$ $A_{tot}$ $Y^{(coll)}_{p}$
@@
Table 1: Summary table of the reactions studied : incident beam energies ($E_{beam}$) and total kinetic energies in the center of mass of the reaction ($K^{(CM)}_{tot}$) are in MeV/nucleon; the velocity of the center of mass of the reaction ($\beta^{(lab)}_{CM}=p^{(lab)}_{tot}/E^{(lab)}_{tot}$) around which the transverse plane is defined; the number of nucleons ($A_{tot}$) and the proton fraction of the system ($Y^{(coll)}_{p}=Z_{tot}/A_{tot}$). Ni+Ni reactions were collected at the GANIL facility with a minimum trigger multiplicity of 4, and the others at the GSI facility with a multiplicity of 3.
@@
For all the reactions, we apply the following method: we select the isotopes emitted in the transverse plane of the reaction around the velocity of the center of mass of the reaction, $\beta^{(lab)}_{CM}$. We define a longitudinal velocity interval $\Delta\beta_{l}=0.02$ and the yields integrated between $\beta_{l}^{-}=\beta^{(lab)}_{CM}-\frac{\Delta\beta_{l}}{2}$ and $\beta_{l}^{+}=\beta^{(lab)}_{CM}+\frac{\Delta\beta_{l}}{2}$ constitute the experimental data for the analysis.
As INDRA is a beam-on-target reaction detector, the direction of the telescopes points to the target position. The angular aperture ($\theta^{(min)}_{r}$,$\theta^{(max)}_{r}$) of the rings must be taken into account in the selection of the transverse plane: $\beta_{l}^{\pm}=\beta^{(lab)}_{CM}\pm\frac{\Delta\beta_{l}}{2}$. If we consider, for a given ring, a particle with velocity ($\beta$), under the isotropic hypothesis, its real longitudinal velocity could be indiscriminately between $\beta^{(min)}_{l}=\beta cos\theta^{(max)}_{r}$ and $\beta^{(max)}_{l}=\beta cos\theta^{(min)}_{r}$. From this uncertainty arises the probability that the particle has been emitted in the transverse plane or not. This is why we apply the following weighting to the production rates detected, $y_{r}^{(det)}(\beta)$: $y_{r}^{(exp)}(\beta)=y_{r}^{(det)}(\beta)\times\frac{\delta^{\cap}_{r}(\beta)}{\delta_{r}}$ with $\delta_{r}=\beta^{(max)}_{l}-\beta^{(min)}_{l}$ and $\delta^{\cap}_{r}(\beta)=\mathrm{min}\left(\beta^{(max)}_{l},\beta^{+}_{l}\right)-\mathrm{max}\left(\beta^{(min)}_{l},\beta^{-}_{l}\right)$. Once this weighting has been applied, the weighted contributions of the various rings are added together and integrated over the velocity range : $y^{(exp)}=\int\sum_{rings}y_{r}^{(exp)}(\beta)\mathrm{d}\beta$. The integrated yields are then normalized to the number of collisions ($N_{coll}$) collected to obtain the mean experimental multiplicities: $m^{(exp)}=\frac{1}{N_{coll}}y^{(exp)}$. These multiplicities will be compared to the prediction of the thermal model.
@@ III.1 Grand Canonical ensemble
with h is the Planck constant, $d=2s+1$ is the spin degeneracy and $\kappa$ indicates which type of statistics is followed by the particle  [^1] If we were in the regim of Fermi-Dirac and Bose-Einstein statistics, $\kappa$=-1 or 1 will be respectively for bosons and fermions. If the particles follow the Maxwell-Boltzmann statistics ($\kappa=0$) and the energy regime allows the single-particle energy to be approximated by $E(p)\sim p^{2}/2M+M$. Then the density number can be written as:
@@ III.2 Two-steps process calculation
This notion of primary distribution comes from the fact that some populated states are unstable. Nucleons populate states that will decay in very short time periods compared to detection times involved in heavy-ion collisions  [^3] The widths of these transient states vary between eV and MeV, and are therefore neglected (set to zero) in view of the masses involved of several GeV. These transient states include both excited states and also the ground state of certain clusters which subsequently referred to as resonances. Their treatment require an additional step in the calculation: after the first step, where the primary distribution of nucleons among neutrons, protons and clusters is computed according to the thermal model equations 2 and 5, a second step is applied: we perform the redistribution of nucleons from the transient states to the stable ground states. In the end only stable states are populated 444the treatment is identical to that performed for resonances in the Resonant Hadron Gas model used at higher energies. The associated density numbers, $n^{(sec.)}$, are written as:
@@ IV.1 Data reproduction
The experimental multiplicities (open black squares) for three of the reactions studied are shown on top panels of figure 1. To these are added the different predictions of the thermal model according to the different cases listed before (full colored symbols). For easier comparison, on bottom panels, we display the difference between prediction and experimental points, $\delta m=m^{(fit)}-m^{(exp)}$. To correctly reproduce the data, the introduction of transient states is necessary. For case 1, where they are not included, we observe discrepancies in the prediction of multiplicities and, specially, the right balance between the 2,3 H and 3 He isotopes is not achieved and this is more pronounced for heavier systems and higher bombarding energies. For the predictions of cases 2 and 3, there are no noticable differences, and we observe a very good agreement with the data. To be quantitative, the NSR estimator (eq. 12) is above 97% for these two latter cases for all reactions included in this comparison while, for case 1, the NSR is around 95% for Ni+Ni reactions and drops to only 90% for Xe+Sn et Au+Au reactions. It should be noted that the 6He multiplicities are slightly less well reproduced. The contribution of transient states, $P_{tr.st.}$, populated during the first step is computed with eq. 13. For case 2, the percentage of nucleons that are redistributed is between 21 and 28% while for case 3, values are between 35 and 45%.
@@ IV.2 Parameter trends
case 2 case 3
Systems $E_{beam}$ T [MeV] $\mu_{B}$ [MeV] $\mu_{Q}$ [MeV] NSR T [MeV] $\mu_{B}$ [MeV] $\mu_{Q}$ [MeV] NSR
58Ni+58Ni 82 6.5 $\pm$ 0.1 927.1 $\pm$ 0.1 -3.5 $\pm$ 0.2 0.99 5.7 $\pm$ 0.1 926.9 $\pm$ 0.2 -3.1 $\pm$ 0.4 0.98
58Ni+58Ni 90 6.7 $\pm$ 0.1 926.8 $\pm$ 0.1 -3.7 $\pm$ 0.2 0.99 5.9 $\pm$ 0.1 926.6 $\pm$ 0.2 -3.2 $\pm$ 0.4 0.98
129Xe+124Sn 80 6.0 $\pm$ 0.3 929.2 $\pm$ 0.3 -7.3 $\pm$ 0.4 0.98 5.3 $\pm$ 0.2 929.0 $\pm$ 0.3 -7.2 $\pm$ 0.4 0.98
124Xe+112Sn 100 6.7 $\pm$ 0.2 927.5 $\pm$ 0.2 -5.8 $\pm$ 0.2 0.99 5.9 $\pm$ 0.0 927.4 $\pm$ 0.1 -5.5 $\pm$ 0.1 0.99
124Xe+124Sn 100 6.7 $\pm$ 0.2 928.0 $\pm$ 0.3 -6.9 $\pm$ 0.3 0.98 5.9 $\pm$ 0.1 927.9 $\pm$ 0.1 -6.8 $\pm$ 0.2 0.99
129Xe+112Sn 100 6.6 $\pm$ 0.2 927.8 $\pm$ 0.2 -6.2 $\pm$ 0.3 0.99 5.9 $\pm$ 0.1 927.6 $\pm$ 0.1 -6.0 $\pm$ 0.1 0.99
129Xe+124Sn 100 6.6 $\pm$ 0.2 928.2 $\pm$ 0.3 -7.3 $\pm$ 0.3 0.98 5.9 $\pm$ 0.1 928.0 $\pm$ 0.1 -7.1 $\pm$ 0.1 0.99
124Xe+124Sn 150 7.9 $\pm$ 0.3 925.7 $\pm$ 0.4 -6.5 $\pm$ 0.5 0.97 6.7 $\pm$ 0.1 925.9 $\pm$ 0.2 -6.3 $\pm$ 0.3 0.98
129Xe+124Sn 150 7.8 $\pm$ 0.3 925.9 $\pm$ 0.4 -6.8 $\pm$ 0.5 0.97 6.7 $\pm$ 0.1 926.2 $\pm$ 0.2 -6.7 $\pm$ 0.3 0.98
197Au+197Au 80 6.1 $\pm$ 0.3 929.4 $\pm$ 0.3 -7.5 $\pm$ 0.4 0.98 5.4 $\pm$ 0.1 929.1 $\pm$ 0.2 -7.5 $\pm$ 0.3 0.98
197Au+197Au 100 6.9 $\pm$ 0.2 928.4 $\pm$ 0.2 -7.4 $\pm$ 0.3 0.98 5.9 $\pm$ 0.3 928.4 $\pm$ 0.4 -7.4 $\pm$ 0.5 0.98
197Au+197Au 150 7.9 $\pm$ 0.3 926.5 $\pm$ 0.3 -6.8 $\pm$ 0.4 0.98 6.7 $\pm$ 0.1 926.6 $\pm$ 0.2 -6.6 $\pm$ 0.3 0.97
Table 2: For each reaction included in the analysis, values and uncertainties of the three thermal model parameters (T, $\mu_{B}$ and $\mu_{Q}$ in MeV) obtain from the fit procedure for the case 2 and 3 (see text for details). The estimator of the reproduction of nucleon sharing (NSR, eq 12) among free nucleons and clusters is also indicated.
@@
Now, if we look in average at the differences in parameter values between case 2 and 3, we mainly observe a negative temperature shift of around $\Delta^{(3-2)}T\sim-0.87$ MeV when excited states are included. For chemical potentials, the values are compatible if we take into account the error bars of the fit ($\Delta^{(3-2)}\mu_{B}\sim-0.08$ and $\Delta^{(3-2)}\mu_{Q}\sim 0.17$ MeV). This systematic difference on temperatures is explained by the opening of accessible states. Indeed, for the same set of parameters, the total density will be greater in case 3 (by around 30% on average). But as the number of nucleons is constrained in the fit (eq. 11), it is the temperature that decreases to counterbalance this opening to excited states without affecting chemical potentials.
@@ IV.3 New constraints on the “freeze-out” line
The right panel of fig. 2 is more focused and information related to the nuclear clustering we can extract from the use of the thermal model. The black symbols and lines are the same as in the left panel. We have also drawn lines framing the points and symbolize the trajectories for two sets of constant values of baryonic density, $n_{B}$, and chemical electric potential, $\mu_{Q}$. These values are the extreme ones obtained from the fit procedure (table 2). For each set of parameters, two lines are drawn illustrating the discussion regarding the inclusion (lower line) or exclusion (upper one) of the excited states in the thermal model. Concerning the density values, 0.008 and 0.020 fm-3, they are consistent with other analysis at lower incident energy [15, 16]. They indicate also that the reproduction of experimental clusters multiplicities with an ideal gas, made of free nucleons and clusters at chemical equilibrium, is associated with low density values of one order of magnitude below the normal density. For comparison, the “freeze-out” line is well reproduced by a constant density line ($n_{B}+n_{\bar{B}}\sim$ 0.12 fm-3) for $\mu_{B}$ values lower than 600 MeV [1]. Concerning the possible comparison of our results with predictions coming from models using baryonic density as main control parameter, the list of particles and available states used to build this density should be known and similar to allow a pertinent comparison.

### v2

@@ II Experimental details
Systems $E_{beam}$ $K^{(c.m.)}_{tot}$ $\beta^{(lab)}_{c.m.}$ $A_{tot}$ $Y^{(coll)}_{p}$
@@
Table 1: Summary table of the reactions studied : incident beam energies ($E_{beam}$) and total kinetic energies in the center of mass of the reaction ($K^{(c.m.)}_{tot}$) are in MeV/nucleon; the velocity of the center of mass of the reaction ($\beta^{(lab)}_{c.m.}=p^{(lab)}_{tot}/E^{(lab)}_{tot}$) around which the transverse plane is defined; the number of nucleons ($A_{tot}$) and the proton fraction of the system ($Y^{(coll)}_{p}=Z_{tot}/A_{tot}$). Ni+Ni reactions were collected at the GANIL facility with a minimum trigger multiplicity of 4, and the others at the GSI facility with a multiplicity of 3.
@@
For all the reactions, we apply the following method: we select the isotopes emitted in the transverse plane of the reaction around the velocity of the center of mass of the reaction, $\beta^{(lab)}_{c.m.}$. We define a longitudinal velocity interval $\Delta\beta_{l}=0.02$ and the yields integrated between $\beta_{l}^{-}=\beta^{(lab)}_{c.m.}-\frac{\Delta\beta_{l}}{2}$ and $\beta_{l}^{+}=\beta^{(lab)}_{c.m.}+\frac{\Delta\beta_{l}}{2}$ constitute the experimental data for the analysis.
As INDRA is a beam-on-target reaction detector, the direction of the telescopes points to the target position. The angular aperture ($\theta^{(min)}_{r}$,$\theta^{(max)}_{r}$) of the rings must be taken into account in the selection of the transverse plane: $\beta_{l}^{\pm}=\beta^{(lab)}_{c.m.}\pm\frac{\Delta\beta_{l}}{2}$. If we consider, for a given ring, a particle with velocity ($\beta$), under the isotropic hypothesis, its real longitudinal velocity could be indiscriminately between $\beta^{(min)}_{l}=\beta cos\theta^{(max)}_{r}$ and $\beta^{(max)}_{l}=\beta cos\theta^{(min)}_{r}$. From this uncertainty arises the probability that the particle has been emitted in the transverse plane or not. This is why we apply the following weighting to the production rates detected, $y_{r}^{(det)}(\beta)$: $y_{r}^{(exp)}(\beta)=y_{r}^{(det)}(\beta)\times\frac{\delta^{\cap}_{r}(\beta)}{\delta_{r}}$ with $\delta_{r}=\beta^{(max)}_{l}-\beta^{(min)}_{l}$ and $\delta^{\cap}_{r}(\beta)=\mathrm{min}\left(\beta^{(max)}_{l},\beta^{+}_{l}\right)-\mathrm{max}\left(\beta^{(min)}_{l},\beta^{-}_{l}\right)$. Once this weighting has been applied, the weighted contributions of the various rings are added together and integrated over the velocity range : $y^{(exp)}=\int\sum_{rings}y_{r}^{(exp)}(\beta)\mathrm{d}\beta$. The integrated yields are then normalized to the number of collisions ($N_{coll}$) collected to obtain the mean experimental multiplicities: $m^{(exp)}=\frac{1}{N_{coll}}y^{(exp)}$. These multiplicities will be compared to the prediction of the thermal model. All the multiplicity values used in the following are listed in [6].
@@ III.1 Grand Canonical ensemble
with h is the Planck constant, $d=2s+1$ is the spin degeneracy and $\kappa$ indicates which type of statistics is followed by the particle  [^1] If we were in the regim of Fermi-Dirac and Bose-Einstein statistics, $\kappa$=-1 or 1 will be respectively for bosons and fermions. If the particles follow the Maxwell-Boltzmann statistics ($\kappa=0$) and the energy regime allows the single-particle energy to be approximated by $E(p)\approx p^{2}/2M+M$. Then the density number can be written as:
@@ III.2 Two-steps process calculation
This notion of primary distribution comes from the fact that some populated states are unstable. Nucleons populate states that will decay in very short time periods compared to detection times involved in heavy-ion collisions  [^3] The widths of these transient states vary between eV and MeV, and are therefore neglected (set to zero) in view of the masses involved of several GeV. These transient states include both excited states and also the ground state of certain clusters which subsequently referred to as resonances. Their treatment require an additional step in the calculation: after the first step, where the primary distribution of nucleons among neutrons, protons and clusters is computed according to the thermal model equations 2 and 5, a second step is applied: we perform the redistribution of nucleons from the transient states to the stable ground states. In the end only stable states are populated 444the treatment is similar to that performed for baryonic resonances in the Resonant Hadron Gas model used at higher energies. The associated density numbers, $n^{(sec.)}$, are written as:
@@ IV.1 Data reproduction
The experimental multiplicities (open black squares) for three of the reactions studied are shown on top panels of figure 1. To these are added the different predictions of the thermal model according to the different cases listed before (full colored symbols). For easier comparison, on bottom panels, we display the difference between prediction and experimental points, $\delta m=m^{(fit)}-m^{(exp)}$. To correctly reproduce the data, the introduction of transient states is necessary. For case 1, where they are not included, we observe discrepancies in the prediction of multiplicities and, specially, the right balance between the 2,3 H and 3 He isotopes is not achieved and this is more pronounced for heavier systems and higher bombarding energies. For the predictions of cases 2 and 3, there are no noticable differences, and we observe a very good agreement with the data. We thus confirm the need to take transient states into account in this type of comparison, as already mentioned in the conclusion of [9]. To be quantitative, the NSR estimator (eq. 12) is above 97% for these two latter cases for all reactions included in this comparison while, for case 1, the NSR is around 95% for Ni+Ni reactions and drops to only 90% for Xe+Sn et Au+Au reactions. It should be noted that the 6He multiplicities are slightly less well reproduced. The contribution of transient states, $P_{tr.st.}$, populated during the first step is computed with eq. 13. For case 2, the percentage of nucleons that are redistributed is between 21 and 28% while for case 3, values are between 35 and 45%.
@@ IV.2 Parameter trends
Reactions case 2 case 3
Systems $E_{beam}$ T [MeV] $\mu_{B}$ [MeV] $\mu_{Q}$ [MeV] $n_{B}$[nuc.fm-3] T [MeV] $\mu_{B}$ [MeV] $\mu_{Q}$ [MeV] $n_{B}$[nuc.fm-3]
58Ni+58Ni 82 6.5$\pm$0.1 927.1$\pm$0.1 -3.5$\pm$0.2 0.015$\pm$0.002 5.7$\pm$0.1 926.9$\pm$0.2 -3.1$\pm$0.4 0.010$\pm$0.003
58Ni+58Ni 90 6.7$\pm$0.1 926.8$\pm$0.1 -3.7$\pm$0.2 0.014$\pm$0.002 5.9$\pm$0.1 926.6$\pm$0.2 -3.2$\pm$0.4 0.010$\pm$0.003
129Xe+124Sn 80 6.0$\pm$0.3 929.2$\pm$0.3 -7.3$\pm$0.4 0.010$\pm$0.004 5.3$\pm$0.2 928.9$\pm$0.3 -7.2$\pm$0.4 0.006$\pm$0.002
124Xe+112Sn 100 6.6$\pm$0.2 927.6$\pm$0.2 -5.8$\pm$0.2 0.012$\pm$0.002 5.9$\pm$0.0 927.4$\pm$0.1 -5.5$\pm$0.1 0.008$\pm$0.001
124Xe+124Sn 100 6.7$\pm$0.3 928.0$\pm$0.3 -6.9$\pm$0.3 0.012$\pm$0.004 5.9$\pm$0.1 927.9$\pm$0.1 -6.8$\pm$0.2 0.007$\pm$0.001
129Xe+112Sn 100 6.6$\pm$0.2 927.8$\pm$0.2 -6.2$\pm$0.3 0.012$\pm$0.003 5.9$\pm$0.1 927.6$\pm$0.1 -6.0$\pm$0.1 0.008$\pm$0.001
129Xe+124Sn 100 6.6$\pm$0.2 928.2$\pm$0.2 -7.3$\pm$0.3 0.011$\pm$0.003 5.9$\pm$0.1 928.0$\pm$0.1 -7.1$\pm$0.1 0.007$\pm$0.001
124Xe+124Sn 150 7.9$\pm$0.5 925.7$\pm$0.6 -6.5$\pm$0.6 0.012$\pm$0.006 6.7$\pm$0.1 925.9$\pm$0.2 -6.3$\pm$0.3 0.008$\pm$0.002
129Xe+124Sn 150 7.8$\pm$0.3 925.9$\pm$0.4 -6.8$\pm$0.5 0.012$\pm$0.004 6.7$\pm$0.1 926.2$\pm$0.2 -6.7$\pm$0.3 0.007$\pm$0.002
197Au+197Au 80 6.1$\pm$0.3 929.4$\pm$0.3 -7.5$\pm$0.4 0.013$\pm$0.005 5.4$\pm$0.1 929.1$\pm$0.2 -7.5$\pm$0.3 0.008$\pm$0.002
197Au+197Au 100 6.9$\pm$0.2 928.4$\pm$0.2 -7.4$\pm$0.3 0.016$\pm$0.004 5.9$\pm$0.3 928.4$\pm$0.4 -7.4$\pm$0.5 0.009$\pm$0.005
197Au+197Au 150 7.9$\pm$0.3 926.5$\pm$0.3 -6.8$\pm$0.4 0.018$\pm$0.005 6.7$\pm$0.1 926.6$\pm$0.2 -6.6$\pm$0.3 0.010$\pm$0.002
Table 2: For each reaction included in the analysis, values and uncertainties of the three thermal model parameters (T, $\mu_{B}$ and $\mu_{Q}$ in MeV) obtain from the fit procedure for the case 2 and 3 (see text for details). Values of deduced baryonic densities ($n_{B}=\sum_{i}A_{i}n_{i}$) are also listed. The uncertainties are obtained by independently exploring the uncertainties of the T, $\mu_{B}$ and $\mu_{Q}$ parameters. The values displayed in figures 2 and 3 are the average of the values for the two cases.
@@
Now, if we look in average at the differences in parameter values between case 2 and 3, we mainly observe a negative temperature shift of around $\Delta^{(3-2)}T\approx-0.87$ MeV when excited states are included. For chemical potentials, the values are compatible if we take into account the error bars of the fit ($\Delta^{(3-2)}\mu_{B}\approx-0.08$ and $\Delta^{(3-2)}\mu_{Q}\approx 0.17$ MeV). This systematic difference on temperatures is explained by the opening of accessible states. Indeed, for the same set of parameters, the total density will be greater in case 3 (by around 30% on average). But as the number of nucleons is constrained in the fit (eq. 11), it is the temperature that decreases to counterbalance this opening to excited states without affecting chemical potentials. A similar shift to lower densities are consequently observed for the case 3.
@@ IV.3 New constraints on the “freeze-out” line
The right panel of fig. 2 is more focused and information related to the nuclear clustering we can extract from the use of the thermal model. The black symbols and lines are the same as in the left panel. We have also drawn lines framing the points and symbolize the trajectories for two sets of constant values of baryonic density, $n_{B}$, and chemical electric potential, $\mu_{Q}$. These values are the extreme ones obtained from the fit procedure (table 2). For each set of parameters, two lines are drawn illustrating the discussion regarding the inclusion (lower line) or exclusion (upper one) of the excited states in the thermal model [^7] For comparison, the freeze-out line is well reproduced by a constant density line ($n_{B}+n_{\bar{B}}\approx$ 0.12 fm-3) for $\mu_{B}$ values lower than 600 MeV [1]..
Now, if we look at the density values deduced from solutions obtained with the thermal model, they indicate that the reproduction of experimental clusters multiplicities with an ideal gas, made of free nucleons and clusters at chemical equilibrium, is associated with low density values of one order of magnitude below the normal density. They are also qualitatively compatible with those obtained in cluster analyses performed on data collected at Fermi energies [11, 12, 13, 14]. There are, however, differences in the computation of density between the two approaches. In our case, this density is obtained by taking into account all the states accessible to nucleons (list given in the subsection III.3) and deduced from the thermal model after ensuring that the data are reproduced and compatible with equilibrium. For the analyses mentioned, densities are obtained using Albergo’s formulation [15] or dedicated derivations. This formulation postulates an a-priori chemical equilibrium to simplify equation 2) and derive density and temperature expressions in terms of isotopic ratios [^8] Albergo’s formulation also makes the following approximation for cluster masses: $M_{i}=A_{i}\times 931.49$, so that only binding energies come into play. However, if we wish to extract chemical potential values, this approximation is not suited.. It also makes the strong assumption of a truncation of the set of accessible states: only free nucleons and stable clusters (generally 2,3H and 3,4He) come into play during the clusterization process, and no de-excitation process of transient states exists or participates in the population of these clusters. In [16], a variable is proposed ($R_{test}$) to quantify the deviation of the data from the case of chemical equilibrium corresponding to $R_{test}$=1. However, this variable is also based on this truncation of accessible states. Indeed, if we calculate $R_{test}$ from equation 10, with the parameter values obtained in this work (table 2), we obtain $R_{test}$ values ranging from 1.29 to 1.38. This confirms that, depending on the accessible states included in the calculation, conclusions about the interpretation of experimental data can change. To sum up, it would be interesting to apply our method to the experimental multiplicities used for these published low-energy analyses, to see whether taking transient states into account can reproduce the data. For the moment, these data are reproduced with assumptions of in-medium effects implemented in various ways in different models [16, 17]. It could be instructive, for example, to relate our $\mu_{Q}$ values to the Coulomb effects mentioned in [13]. Concerning the possible comparison of our results with predictions coming from models using baryonic density as main control parameter, the list of particles and available states used to build this density should be known and similar to allow a pertinent comparison.